# Understanding Differences in News Article Interaction Patterns on Facebook: Public vs. Private Sharing with Varying Bias and Reliability

### v2

@@ 3 Dataset
4 Public vs. Private Interactions
To better understand the differences in interactions with public vs. private posts, we first compare the interaction rates with all public posts and with all combined (public + private) posts. We note that this implicit comparison approach provides more stable and reliable statistical comparisons than trying to use only the estimated number of private interactions. The relative strength of this approach is further compounded by the private interactions typically dominating the combined set.
4.1 Aggregated (Macro) Analysis of Interactions
Figure 3: Normalized number of combined and public interactions per article for each bias and reliability class.
@@ 4.1 Aggregated (Macro) Analysis of Interactions
  • Average combined interactions (left bars for each class) represents the average number of combined (private+public) interactions per article in each class. More specifically, considering $I_{c,i}^{\text{comb}}$ as the combined interaction for the $i^{th}$ article in class $c$, the average combined interactions for this class ($\bar{I}_{c}^{\text{comb}}$) is calculated as:
    $\bar{I}_{c}^{\text{comb}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}{N_{c}},$ (1)
@@
  • Average public interactions per article (right bars) is defined in a similar manner but considering only the public interactions. More specifically, considering $I_{c,i}^{\text{public}}$ as the public interactions for the $i^{th}$ article in a class $c$, the average public interactions for this class ($\bar{I}_{c}^{\text{public}}$) is calculated as follows:
    $\bar{I}_{c}^{\text{public}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{N_{c}},$ (2)
@@
As seen in the figure, there are some very clear and interesting differences in the trends observed for the public vs. the combined statistics. First, for all three bias classes and for all four reliability classes, the interaction rates are significantly lower for the public posts than for the combined set (which for all classes are dominated by private posts). These aggregate rate differences suggest that users are more likely to interact with private posts, regardless of bias and reliability class, and underscores the importance of considering the total (combined) interactions, not only the (typically studied) public posts. We next look closer at the relative interaction levels of each class and the impact of using the public vs. combined sets for such interaction comparisons.
Bias comparisons: Second, looking at the relative differences among the bias classes, the Left bias class has the highest average combined interactions per article (17,501), followed by the Right (16,325) and Center (8,120) classes. For public interactions; however, the Right bias class has the highest average public interactions per article (5,236), with the Left (4,497) and Center (2,894) classes trailing behind. This shows that biased articles see even greater relative interaction rates when posts are private than public.
Reliability comparisons: Third, for the reliability classes, the Most-reliable class has the highest average combined and public interactions per article (17,995 and 5,101). The Unreliable and Reliable classes have relatively similar average public interactions per article (3,741 and 3,291, respectively). Moreover, the Most-unreliable class has a higher average combined interactions per article (12,869) and public interactions per article (3,751). It is interesting to see that the two extreme classes (i.e., the Most-reliable and Most-unreliable) have the highest average interactions per article in both combined and public contexts, implying that users are more likely to engage with articles on both ends of the reliability spectrum, regardless if the posts are public or non-public, although the level of interactions is different.
Key observation: Compared to the Center group, biased groups receive higher levels of aggregated engagement per article. In terms of reliability, the Reliable news reaches the lowest engagement levels per article.
Aggregated public interaction share: Motivated by most prior works only studying the public posts tracked by CrowdTangle, we next look closer at how big fraction of the total interactions this set captures and compare the relative differences in the public interaction share of the different categories. For this analysis, we calculate the (aggregate) public interaction share, defined as the ratio of public interactions to the combined interactions for each class $c$ as $\bar{I}_{c}^{\text{public}}/\bar{I}_{c}^{\text{comb}}$.
Figure 4: Aggregated public interactions shares
@@
Key observation: Compared to right-biased articles, left-biased articles receive higher private interactions share. In terms of reliability, extremely reliable/unreliable news articles receive higher levels of private interactions share.
4.2 Granular Statistical (Micro) Analysis of Interactions
@@
The next two subsections compare the distributions of the total (combined) interactions (Sect. 4.2.1) and study the distributions of the relative public interaction shares (Sect. 4.2.2), respectively. In both subsections, the distributions are compared using multiple statistical tests; statistical significance is reported if a p-value is below 0.001, and supporting p-values are reported for the main findings.
4.2.1 Distribution of Combined Interactions
[…]
Figure 5: Empirical complementary cumulative distribution function (CCDF) of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different bias classes. While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
@@ 4.2.1 Distribution of Combined Interactions
Key observation: Statistical analysis further strengthens the observation that biased news consistently garners higher combined engagement levels per article than center-aligned news. Interestingly, there is no statistically significant difference between the Left and Right class.
@@
Comparison of Reliability Classes: Figure 6 shows the CCDFs for the different reliability classes, broken down for both combined and public interactions, accompanied by a boxplot offering a clearer perspective on the percentile values for the combined interactions. The CCDFs again demonstrate heavy-tailed distribution characteristic, and the relative shifts of the distributions are consistent with the aggregate (average values) previously discussed (and seen in Figure 3), with the Most-reliable class typically getting the most interactions, followed by the unreliable classes, and subsequently by the Reliable class. While comparing the tail parts is straightforward, comparing the whole distributions requires some care. For example, here, the core of the boxplot does not distinctly set apart the Most-reliable from the unreliable classes. To derive robust conclusions, we therefore again performed a sequence of statistical tests on the combined interactions distributions.
Figure 6: CCDF of both combined and public interactions, alongside an inset boxplot detailing the distribution of the combined interactions for different reliability classes. While the CCDF underscores the tail characteristics of the distributions, the boxplot provides a granular statistical summary of the combined interactions, including medians, quartiles, and means.
@@
First, the Kruskal-Wallis test was applied to discern differences among the reliability classes. While we observed statistical differences (p-value of $7.9\cdot 10^{-20}$), the magnitude of distinction among reliability classes is more nuanced than in the bias classes (which had a p-value of $1.66\cdot 10^{-65}$). Further analysis employing the Dunn’s test confirmed that the Reliable class statistically diverges from the other three. The least significant p-value here, $7.63\cdot 10^{-8}$, is attributed to the comparison between the Reliable and Unreliable classes. The most significant, on the other hand, emerges from the Reliable versus Most-reliable comparison with the p-value of $7.05\cdot 10^{-15}$. Except for these distributions, none of the other pairwise distribution comparisons resulted in statistically significant differences. Third, the Mann-Whitney U test reinforced our findings, particularly highlighting the statistical significance when comparing the medians of the Reliable class (median of 891) with the others: Most-unreliable (1,162), Unreliable (1,104), and Most-reliable (1,218). While all three pairwise comparisons are significant, their p-values range from $8.96\cdot 10^{-9}$ (Reliable vs. Unreliable) to $8.4\cdot 10^{-16}$ (Reliable vs. Most-reliable). Finally, we apply bootstrapping to compare the means. While only the Reliable and Most-reliable classes comparison is significant at the 99% confidence level, we note that all three comparisons against the Reliable class are significant at the 95% confidence level.
@@
Key observation: Comparing to the other reliability classes, the Reliable articles demonstrate statistically significant lower interaction values, especially when compared to the Most-reliable articles.
4.2.2 Distribution of Public Interaction Shares
@@
Figure 7: Distribution of public interaction ratios
@@
Key observation: The two extreme classes, Most-reliable and Most-unreliable, exhibit similar private interaction ratios that are statistically higher compared to the other classes.
We now shift our attention to the public interaction ratios across the bias classes, where we make some interesting observations regarding the classes’ relative order. Specifically, we now observe the first difference in the relative order of the classes when comparing the average (and median) public shares calculated on a per-article basis (Figure 7) with the aggregated results (Figure 4). More specifically, comparing with the aggregate results, the Right class has now switched ranking with the Left class to become the class with the noticeably smallest public interaction share. Furthermore, comparing the distributions seen for the Right and Left class is statistically significant: Dunn test (p-value of $3.25\times 10^{-14}$), Mann-Whitney U test (p-value of $6.16\times 10^{-17}$) and bootstrap results being significant at 99% level.
While these results may appear surprising at first, they are due to some interesting differences in the distributions. First, note that the aggregate analysis only calculates the overall ratio $\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}$ compared to the granular analysis, where the average (for example) is calculated as $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\frac{I_{c,i}^{\text{public}}}{I_{c,i}^{\text{comb}}}$. Second, taking a closer look at the distributions, $\mathcal{D}_{Left}^{PIR}$ and $\mathcal{D}_{Right}^{PIR}$ differ substantially. Part of this can be observed in Figure 5, where we can see that Left class demonstrates notable disparities between public and combined interactions in the tail, while the Right class exhibits relatively larger differences in the head of the distribution. This suggests that articles with higher combined number of interactions in the Right class may exhibit higher normalized public interaction shares, and articles with few combined interactions may exhibit relatively smaller public shares (compared to the Left class).
@@
Key observation: In terms of bias, the Right class sees the smallest public interaction share among the least popular content to interact with, while the Left class sees relatively less overall public interaction share (due to its most popular articles to interact with not seeing as big public interaction share).
5 Deep vs. Shallow Interactions
In this section, we investigate the prevalence of deep vs. shallow interactions and explore their variations among different reliability and bias classes. For this analysis, we categorize all (emoji-based) reactions (i.e., likes, loves, sads, etc.) as shallow interactions while comments and shares are considered as deep interactions. This distinction, supported by the previous works [8, 33], is motivated by the idea that comments and shares typically involve more cognitive effort and engagement from users, as they require formulating thoughts, opinions, or responses in the form of comments or actively endorsing and disseminating the content by sharing. In contrast, reacting with an emoji is a more passive and less cognitively demanding form of interaction.
@@
For example, consider an article with 1,000 total interactions, 500 of which are deep interactions, resulting in a DIRC of 0.5. Now, if 400 of these interactions are public, and 100 of them are deep interactions, the DIRP is 0.25 for the same article. We next analyze the distribution of DIRC and DIRP values for all articles in each class.
@@
Figure 8: Deep interaction ratios (DIRC and DIRP) for different bias classes and a baseline.
First, referring to the figure, for each bias class, we observe substantial difference between the distributions of DIRC and DIRP. For example, for every percentile and for the means, the DRIC distributions (combined) have significantly larger values than the DIRP distributions (public). This again highlights the difference between the public and combined interactions and the importance of considering both the private and public spheres in future research. This observation leads us to the following key insight.
Key observation: Irrespective of the news class, users tend to engage more deeply in private discussions compared to public ones.
Now, and perhaps even more importantly, to better understand to what extent the relative level of interaction seen for different bias classes depends on which dataset was used we apply the same statistical testing methodology as discussed in previous sections for the two sets independently. First, the Kruskal-Wallis test provides evidence that the distributions differ significantly among the bias classes, regardless if using of the DIRC and DIRP, with p-values of $7.14\cdot 10^{-23}$ and $2.08\cdot 10^{-9}$, respectively. Comparing the p-values, we note that the differences appear more significant using the combined set (DIRC) than using only the public data (DIRP). Second, consistent with visual comparison in the figure, the post-hoc Dunn test only supports the statistically significance differences between the Right group (which has the biggest fraction of deep interactions regardless of using the combined or pubic sets) and the other two groups but not among the Center and Left group. Finally, the Mann-Whitney U test and bootstrapping support the same pattern for the median and means, with the only significant differences again being between the Right class and the other two bias classes. For the bias classes, the main results regarding the relative depth of the interactions among the classes are, therefore, consistent regardless of the dataset.
Figure 9: CDFs of DIRC (combined) and DIRP (public) for different bias classes.
@@
Key observation: Facebook users tend to engage more deeply with content from the Right bias class in both public and combined contexts.
@@
Comparison of reliability classes: We now shift our focus to the reliability classes and their deep interaction ratios. Figure 10 summarizes these results. While we again observe deeper interactions for the combined set (suggesting shallower interactions with public posts), the relative differences among the reliability classes themselves (given a dataset) are less apparent compared to those observed among the bias classes. We also observe some smaller differences in which relative differences are significant, highlighting the value of also considering the combined data (not only the public data).
Figure 10: Deep interaction ratios (DIRC and DIRP) for different reliability classes
Figure 11: CDFs of DIRC (combined) and DIRP (public) for different reliability classes.
@@
First, starting with the combined dataset (DIRC), we observe presence of significant differences among the distributions (Kruskal-Wallis test), with the post-hoc Dunn test (distribution) and Mann-Witney U test (median) revealing that only the Most-unreliable class differs statistically from the others. This observation is further visually supported by the CDF plot in Figure 11, which clearly shows a shift to the right in the DIRC distribution of the Most-unreliable class.
@@
Switching our focus to the DIRP distributions, containing only public interactions, we interestingly observe a slightly different pattern, as we, for that dataset, also observe statistical differences between some additional classes. First, as suggested by the DIRP CDFs in Figure 11, both extreme classes (Most-unreliable and Most-reliable) are statistically different from the other two classes (Kruskal-Wallis test followed by Mann-Whitney test, followed by the post-hoc Dunn test). Second, no statistical significance was observed between the Most-reliable and Most-unreliable classes themselves or between the Reliable and Unreliable classes. We base the following insight based on our observations across the two datasets.

### v3

@@ 3. Dataset
4. Public vs. Combined Interactions
To understand how engagement patterns differ between public and private spheres on Facebook, we present two complementary analyses: (1) an aggregated comparison of public vs. combined (public + private) interactions (Sec. 4.1), and (2) a statistical analysis of their differences (Sec. 4.2). By comparing public interactions against combined interactions, where private interactions constitute the majority (as we show later), we can effectively study how engagement in private spaces differs from public ones.
4.1. Aggregated (Macro) Analysis of Interactions
@@
Average combined interactions: (left bars for each class) represents the average number of combined (private+public) interactions per article in each class. More specifically, considering $I_{c,i}^{\text{comb}}$ as the combined interaction for the $i^{th}$ article in class $c$, the average combined interactions for this class ($\bar{I}_{c}^{\text{comb}}$) is calculated as:
(1) $\bar{I}_{c}^{\text{comb}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}{N_{c}},$
@@
Average public interactions per article: (right bars) is defined in a similar manner but considering only the public interactions. More specifically, considering $I_{c,i}^{\text{public}}$ as the public interactions for the $i^{th}$ article in a class $c$, the average public interactions for this class ($\bar{I}_{c}^{\text{public}}$) is calculated as follows:
(2) $\bar{I}_{c}^{\text{public}}=\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{N_{c}},$
@@
As seen in the figure, there are some very clear and interesting differences in the trends observed for the public vs. the combined statistics. First, for all three bias classes and for all four reliability classes, the interaction rates are significantly lower for the public posts than for the combined set. These aggregate rate differences suggest that users are more likely to interact with private posts, regardless of bias and reliability class, and underscores the importance of considering the total (combined) interactions, not only the (typically studied) public posts. We next look closer at the relative interaction levels of each class and the impact of using the public vs. combined sets for such interaction comparisons.
Bias comparisons: Second, examining the relative differences among the bias classes, the Left bias class has the highest average combined interactions per article (17,501), followed by the Right (16,325) and Center (8,120) classes. For public interactions; however, the Right bias class has the highest average public interactions per article (5,236), with the Left (4,497) and Center (2,894) classes trailing behind. This shows that biased articles see even greater relative interaction rates in private than in public posts.
Reliability comparisons: Third, for the reliability classes, the Most-reliable class has the highest average combined and public interactions per article (17,995 and 5,101). The Unreliable and Reliable classes have relatively similar average public interactions per article (3,741 and 3,291, respectively). Moreover, the Most-unreliable class has a higher average combined interactions per article (12,869) and public interactions per article (3,751). It is interesting to see that the two extreme classes (i.e., the Most-reliable and Most-unreliable) have the highest average interactions per article in both combined and public contexts, implying that users are more likely to engage with articles on both ends of the reliability spectrum.
Aggregated public interaction share: Motivated by most prior works only studying the public posts tracked by CrowdTangle, we next evaluate what fraction of the total interactions this set captures and compare the relative differences in the public interaction share of different categories. For this analysis, we calculate the (aggregate) public interaction share as the ratio of public interactions to total interactions for each class $c$: $\bar{I}_{c}^{\text{public}}/\bar{I}_{c}^{\text{comb}}$.
@@
Takeaway: Compared to right-biased articles, left-biased articles receive higher private interactions share. In terms of reliability, extremely reliable/unreliable news articles receive higher levels of private interactions share.
Figure 3. Aggregated public interactions shares
(a) Disributions of different bias classes.
(b) Disributions of different reliability classes.
Figure 4. Empirical complementary cumulative distribution functions (CCDFs) for combined and public interactions, along with inset boxplots showing the distribution of combined interactions. Both CCDFs emphasize tail characteristics of distributions, while the boxplots provide granular statistical summaries, including medians, quartiles, and means.
4.2. Granular Statistical (Micro) Analysis of Interactions
@@
The next two subsections compare distributions of the total (combined) interactions (Sect. 4.2.1) and study the distributions of the relative public interaction shares (Sect. 4.2.2), respectively. In both subsections, the distributions are compared using multiple statistical tests; statistical significance is reported if a p-value is below 0.001, and supporting p-values are reported for the main findings.
4.2.1. Distribution of Combined Interactions
Comparison of Bias Classes: Figure 4(a) shows the CCDFs of both the set of combined interactions ($\mathcal{D}^{\text{comb}}_{c}$) and the set of public interactions ($\mathcal{D}^{\text{public}}_{c}$) for each bias class $c$ on a log-log scale. Accompanying the CCDF is a boxplot, highlighting key statistical percentiles for the distribution of the combined interactions ($\mathcal{D}^{\text{comb}}_{c}$). Specifically, the 10th percentile (bottom marker), 25th percentile (bottom of box), median (middle marker), 75th percentile (top of box), 90th percentile (top marker), and the mean (i.e., $\bar{I}_{c}^{\text{comb}}$) (circle). To allow higher resolution, outliers are not shown in the boxplot.
@@ 4.2.1. Distribution of Combined Interactions
Takeaway: Statistical analysis further strengthens the observation that biased news consistently garners higher combined engagement levels per article than center-aligned news. Interestingly, there is no statistically significant difference between the Left and Right class.
@@
Comparison of Reliability Classes: Figure 4(b) shows the CCDFs for the different reliability classes, broken down for both combined and public interactions, accompanied by a boxplot offering a clearer perspective on the percentile values for the combined interactions. The CCDFs again demonstrate heavy-tailed distribution characteristic, and the relative shifts of the distributions are consistent with the aggregate (average values) previously discussed (and seen in Figure 2), with the Most-reliable class typically getting the most interactions, followed by the unreliable classes, and subsequently by the Reliable class. While comparing the tail parts is straightforward, comparing the whole distributions requires some care. For example, here, the core of the boxplot does not distinctly set apart the Most-reliable from the unreliable classes. To derive robust conclusions, we, therefore, again performed a sequence of statistical tests on the combined interactions distributions.
@@
First, the Kruskal-Wallis test was applied to discern differences among the reliability classes. While we observed statistical differences (p-value of $7.9\cdot 10^{-20}$), the magnitude of distinction among reliability classes is more nuanced than in the bias classes (which had a p-value of $1.66\cdot 10^{-65}$). Further analysis employing the Dunn’s test confirmed that the Reliable class statistically diverges from the other three. The least significant p-value here, $7.63\cdot 10^{-8}$, is attributed to the comparison between the Reliable and Unreliable classes. The most significant, on the other hand, emerges from the Reliable versus Most-reliable comparison with the p-value of $7.05\cdot 10^{-15}$. Except for these distributions, none of the other pairwise distribution comparisons resulted in statistically significant differences. Third, the Mann-Whitney U test reinforced our findings, particularly highlighting the statistical significance when comparing the medians of the Reliable class (median of 891) with the others: Most-unreliable (1,162), Unreliable (1,104), and Most-reliable (1,218). Finally, we apply bootstrapping to compare the means. While only the Reliable and Most-reliable classes comparison is significant at the 99% confidence level, we note that all three comparisons against the Reliable class are significant at the 95% confidence level.
@@
Takeaway: Comparing to the other reliability classes, the Reliable articles demonstrate lower interaction values, especially when compared to the Most-reliable articles.
4.2.2. Distribution of Public Interaction Shares
@@
Figure 5. Distribution of public interaction ratios
Figure 5 presents the public interaction ratios using box-plots broken down per bias and reliability category. Here, we again show the 10th percentile, 25th percentile, median, 75th percentile, and 90th percentile, as well as the average. Like for the aggregate analysis, we observe some notable differences, which we can now support using statistical analysis. First, the Kruskal-Wallis test shows that there are statistically significant differences between both the biases classes (p-value of $2.68\cdot 10^{-22}$) and reliability classes (p-value of $8.74\cdot 10^{-33}$).
Takeaway: The two extreme classes, Most-reliable and Most-unreliable, exhibit similar private interaction ratios that are statistically higher compared to the other classes.
We now shift our attention to the public interaction ratios across the bias classes, where we make some interesting observations regarding the classes’ relative order. Specifically, we now observe the first difference in the relative order of the classes when comparing the average (and median) public shares calculated on a per-article basis (Figure 5) with the aggregated results (Figure 3). More specifically, comparing with the aggregate results, the Right class has switched ranking with the Left class, becoming the one with smallest public interaction share. Furthermore, comparing the distributions seen for the Right and Left class is statistically significant: Dunn test (p-value of $3.25\times 10^{-14}$), Mann-Whitney U test (p-value of $6.16\times 10^{-17}$) and bootstrap results being significant at 99% level.
Figure 6. Deep interaction ratios (DIRC and DIRP) for different bias classes and a baseline.
Figure 7. CDFs of DIRC (combined) and DIRP (public) for different bias classes.
@@
Figure 8. CDFs of DIRC (combined) and DIRP (public) for different reliability classes.
[…]
While these results may appear surprising at first, they are due to some interesting differences in the distributions. First, note that the aggregate analysis only calculates the overall ratio $\frac{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{public}}}{\sum_{i=1}^{N_{c}}I_{c,i}^{\text{comb}}}$ compared to the granular analysis, where the average (for example) is calculated as $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\frac{I_{c,i}^{\text{public}}}{I_{c,i}^{\text{comb}}}$. Second, taking a closer look at the distributions, $\mathcal{D}_{Left}^{PIR}$ and $\mathcal{D}_{Right}^{PIR}$ differ substantially. Part of this can be observed in Figure 4(a), where we can see that Left class demonstrates notable disparities between public and combined interactions in the tail, while the Right class exhibits relatively larger differences in the head of the distribution. This suggests that articles with higher combined number of interactions in the Right class may exhibit higher normalized public interaction shares, and articles with few combined interactions may exhibit relatively smaller public shares (compared to the Left class).
@@
Takeaway: In terms of bias, the Right class sees the smallest public interaction share among the least popular content to interact with, while the Left class sees relatively less overall public interaction share (due to its most popular articles to interact with not seeing as big public interaction share).
5. Deep vs. Shallow Interactions
In this section, we investigate the prevalence of deep vs. shallow interactions and explore their variations among different reliability and bias classes. For this analysis, we categorize all (emoji-based) reactions (e.g., likes, loves, sads) as shallow interactions while comments and shares are considered as deep interactions. This distinction, supported by the previous works (Aldous et al., 2019; Kim and Yang, 2017), is motivated by the idea that comments and shares typically involve more cognitive effort and engagement from users. For example, comments involve formulating thoughts and opinions, while shares actively endorse and disseminate content.
@@
To compute DIRC, we divide the number of deep interactions by the total interactions for each URL. For DIRP, we first identify the deep interactions (shares and comments) occurring on public posts associated with the URL. Summing these deep interactions across all public posts and dividing by the total number of public posts’ interactions for the URL yields the DIRP. For example, an article with 1,000 total interactions, 500 of which are deep, has a DIRC of 0.5. If 400 interactions are public, and 100 of those are deep, the DIRP is 0.25. We next analyze the distribution of DIRC and DIRP values for all articles in each class.
@@
First, referring to the figure, for each bias class, we observe substantial differences between the distributions of DIRC and DIRP. For example, for every percentile and for the means, the DIRC distributions (combined) have significantly larger values than the DIRP distributions (public), indicating that users engage in more substantive interactions in less-public settings. This suggests that deeper engagement—through comments and shares—tends to comparatively have higher ratios in spaces with fewer visibility constraints, where users may feel more comfortable expressing opinions or endorsing content. This pattern is consistent across all bias (Figure 8) and reliability classes (Figure 8), reinforcing the notion that private interactions are an essential component of online engagement dynamics. This again highlights the difference between the public and combined interactions and the importance of considering both the private and public spheres in future research. This observation leads us to the following key insight.
Takeaway: Irrespective of the news class, users tend to engage more deeply in private discussions compared to public ones.
Now, and perhaps even more importantly, to better understand to what extent the relative level of interaction seen for different bias classes depends on which dataset was used we apply the same statistical testing methodology as discussed in previous sections for the two sets independently. First, the Kruskal-Wallis test provides evidence that the distributions differ significantly among the bias classes, regardless of using the DIRC and DIRP, with p-values of $7.14\cdot 10^{-23}$ and $2.08\cdot 10^{-9}$, respectively. Comparing the p-values, we note that the differences appear more significant using the combined set (DIRC) than using only the public data (DIRP). Second, consistent with visual comparison in the figure, the post-hoc Dunn test only supports statistically significant differences between the Right group (which has the biggest fraction of deep interactions regardless of using the combined or pubic sets) and the other two groups but not among the Center and Left group. Finally, the Mann-Whitney U test and bootstrapping support the same pattern for the median and means, with the only significant differences again being between the Right class and the other two bias classes. For the bias classes, the main results regarding the relative depth of the interactions among the classes are, therefore, consistent regardless of the dataset.
@@
Takeaway: Content from the Right bias class in both public and combined contexts gets deeper user interactions.
@@
Comparison of reliability classes: We now shift our focus to the reliability classes and their deep interaction ratios. Figure 8 summarizes these results. While we again observe deeper interactions for the combined set (suggesting shallower interactions with public posts), the relative differences among the reliability classes themselves are less apparent compared to those observed among the bias classes. We also observe some smaller differences in which relative differences are significant, highlighting the value of also considering the combined data (not only the public data).
@@
First, starting with the combined dataset (DIRC), we observe the presence of significant differences among the distributions (Kruskal-Wallis test), with the post-hoc Dunn test (distribution) and Mann-Witney U test (median) revealing that only the Most-unreliable class differs statistically from the others. This observation is further visually supported in Figure 8, which clearly shows a shift to the right in the DIRC distribution of the Most-unreliable class.
@@
Switching our focus to the DIRP distributions, containing only public interactions, we interestingly observe a slightly different pattern, as we here also observe statistical differences between some additional classes. First, as suggested by the DIRP CDFs in Figure 8, both extreme classes (Most-unreliable and Most-reliable) are statistically different from the other two classes (Kruskal-Wallis test followed by Mann-Whitney test, followed by the post-hoc Dunn test). Second, no statistical significance was observed between the Most-reliable and Most-unreliable classes themselves or between the Reliable and Unreliable classes. We base the following insight based on our observations across the two datasets.